Solve the following equation:
x-3 = sqrt(x-1)

Answer by edjones(7569)

x-3=sqrt(x-1)
x^2-6x+9=x-1 square each side
x^2-7x+10=0
(x-5)(x-2)=0
x=5, x=2
x=5 is the only answer.
.
Ed